which among log (base 13)169 and log (base 23)521 is greater ?(with out using calculator)
thanks

Answer by lwsshak3(11628)

which among log (base 13)169 and log (base 23)521 is greater
Using exponential form of logs:
Base raised to log of number=number
for first case
13^x=169
x=2 which is the log of 169 base 13
..
for 2nd case
23^x=521
x=<2 which is the log of 529 base 23 (if x=2, the number=23^2=529, so x<2)
The log of the first case>log of 2nd case